# A localized consensus-based sampling algorithm

### v2

@@ Abstract
We develop a novel interacting-particle method for sampling from non-Gaussian distributions. As a first step, we propose a new way to derive the consensus-based sampling (CBS) algorithm, starting from ensemble-preconditioned Langevin diffusions. We approximate the target potential by its Moreau envelope, such that the gradient in the Langevin equation can be replaced by a proximal operator. We then approximate the proximal operator by a weighted mean, and finally assume that the initial and target distributions are Gaussian, resulting in the CBS dynamics. If we keep only those approximations that can be justified in the non-Gaussian setting, the result is a new interacting-particle method for sampling, which we call localized consensus-based sampling. We prove that our algorithm is affine-invariant and exact for Gaussian distributions in the mean-field setting. Numerical tests illustrate that localized CBS compares favorably to alternative methods in terms of affine-invariance and performance on non-Gaussian distributions.
Keywords: Consensus-based sampling $\cdot$ Langevin diffusion $\cdot$ Bayesian inverse problems
@@ 1 Introduction
We consider the problem of sampling from a probability distribution in $d$ dimensions with density
@@
where $V\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a potential function with $\int\exp(-V(v))\mathrm{d}v<\infty$. Here, we implicitly take integrals over $\mathbb{R}^{d}$. In many applications $V$ is known but the integral $Z$ may be intractable, such that $\pi$ is available only up to a multiplicative constant.
@@ 1.1 Bayesian inverse problems
A common application area to give rise to problem Eq. 1 is that of Bayesian inverse problems (see, e.g., [30, 55]). Consider a known model $G$ with unknown parameters $u$ and a perturbed observation
@@
where $\eta$ is some unknown measurement noise. The goal is to characterize likely values of $u$, based on $y$ and $G$. If we denote by the probability densities $\pi_{\mathrm{prior}}(u)$ and $\pi_{\mathrm{noise}}(\eta)$ our prior beliefs about $u$ and $\eta$, then Bayes’ rule describes our belief about $u$ after the observation by the posterior distribution with density
@@
Optimization algorithms can solve the inverse problem in this fashion. However, $u_{\mathrm{MAP}}$ gives no information about the uncertainty on $u$ captured by $\pi_{\mathrm{post}}$. When uncertainty quantification is required, other characteristics of the posterior must be inferred.
In that case, one commonly wants to integrate some function with respect to $\pi_{\mathrm{post}}$, e.g., to compute statistical moments of the distribution. Numerical integration techniques, such as [25], which uses quasi-Monte Carlo, may provide a sufficient approximation even in high dimensions. Alternatively, it is common to use algorithms that produce a sequence of Monte Carlo samples distributed approximately according to $\pi_{\mathrm{post}}$. Then solving Eq. 2 corresponds to drawing samples from Eq. 1, with the evidence $\pi_{\mathrm{ev}}(y)$ as the intractable $Z$ and with $\widehat{\pi}_{\mathrm{post}}(\cdot\mid y)$ as the unnormalized density $\widehat{\pi}$.
@@ 1.2 Related literature
Sampling from distributions with density Eq. 1 with unknown $Z$ is historically performed with Markov chain Monte Carlo (MCMC) algorithms, which originate in [24, 37] and are reviewed in detail in [3]. MCMC remains a very active research topic with recent developments such as the popular pCN [8] and DILI [9] proposal moves that are effective in high dimensions and multilevel MCMC versions [15, 35] that exploit cheap approximations to $G$.
An alternative method is sequential Monte Carlo (SMC) [11], which evolves samples from an easy-to-sample distribution (e.g., $\pi_{\mathrm{prior}}$ in Eq. 3) into samples from $\pi$ (e.g., $\pi_{\mathrm{post}}$) gradually over multiple (time) steps. Many other methods are based on evolving SDEs that have $\pi$ (or a distribution approximating $\pi$) as an invariant distribution. An example is Langevin diffusion (see subsection 2.1); after time discretization, the unadjusted Langevin algorithm (ULA) [49] arises as an approximate sampler of $\pi$. Using ULA steps as MCMC proposals—or, equivalently, adding Metropolis–Hastings rejection to ULA—results in the Metropolis-adjusted Langevin algorithm (MALA) [2].
These Langevin algorithms use derivatives of the potential $V$, which is often undesirable [16, 32]. McKean–Vlasov-type SDEs (SDEs whose dynamics include their own solution law) discretized with multiple particles have been used to construct various derivative-free (often approximate) samplers. Ensemble Kalman sampling (EKS) [19] and the closely related ALDI method [20] are derived from interacting Langevin diffusions and inspired by the ensemble Kalman inversion (EKI) optimizer [28, 32, 52], which is related to a sequential application of the ensemble Kalman filter (EnKF) [17]. There is also a version of EKI that performs sampling, which is analyzed in [13]. The consensus-based sampling (CBS) method [6], described in more detail in subsection 2.2, was inspired by consensus-based optimization (CBO) [46]. Other samplers based on interacting particles include the multiscale sampler in [43].
Interaction in particle methods can be localized: instead of computing a global interaction term, each particle interacts more with particles that are nearby than with those that are far away. The aim of localization is to avoid the linearity assumption on $G$ that is implicit in these methods. Localized algorithms show improved performance for nonlinear $G$, since the adapted dynamics rely on local instead of global linearity. The paper [47] proposed to use localized interaction in the EKS/ALDI algorithm—details are given in subsection 5.3. Variants on this idea were proposed for EKI and the EnKF [59], for an ensemble Kalman method for rare-event sampling [60], and for CBS as a variant named polarized CBS [4]—see subsection 5.2. Our paper develops a different CBS variant with localized interaction, named localized CBS.
@@
We stress that localization as discussed here is unrelated to the notion of localization in [26, 42] for the EnKF, in [1, 58] for EKI, and in [39] for MCMC. There, the empirical covariance matrix is corrected for spurious correlations, caused by a finite ensemble size, between parameter indices that should be uncorrelated—for example, because they correspond to elements at a large distance in the underlying problem. This technique can also increase the rank of the empirical covariance and prevent EKI’s subspace property [58].
@@ 1.3 Objectives and contributions
    Parallelizability. Standard MCMC methods are inherently sequential due to the need to overcome a burn-in phase and are hard to parallelize effectively [54, 61]. Alternatives such as SMC and McKean–Vlasov-based methods use an ensemble of particles in parameter space that exchange information through interaction. This means that any work that is local to a single particle, such as evaluating $G$, can be trivially parallelized over all particles.
@@
    Non-Gaussian performance. Many sampling methods use approximations that hold only for Gaussian distributions, leading to poor sampling performance on non-Gaussian and multimodal target distributions.
@@
    Section 3 proposes a novel way to connect interacting Langevin diffusions [19], which use gradients, to consensus-based sampling (CBS) [6], which does not. This connection is based on three approximations: replacing $V$ by a Moreau envelope, approximating the proximal operator by a weighted mean, and assuming that the initial and target distributions are Gaussian.
@@
    Subsections 4.1 and 4.2 adapt these approximations, removing those that go against our objective of solving non-Gaussian sampling problems. This results in an algorithm we name localized consensus-based sampling. Section 5 compares the existing CBS, polarized CBS, and localized ALDI algorithms to our localized CBS method.
@@
    We perform an analysis of localized CBS with a broad class of covariance-based preconditioners. We derive ODEs for the evolution of its mean and covariance when applied to Gaussian problems in subsection 4.3. More generally, subsection 4.4 shows that localized CBS is affine-invariant.
@@
    Section 6 discusses how to efficiently approximate the mean-field equations with finite particle systems, along with other computational considerations. Section 7 performs numerical experiments that showcase the affine-invariance and non-Gaussian performance of localized CBS.
@@ 1.4 Notation
We define the norm of a vector $v\in\mathbb{R}^{n}$ weighted by a positive definite matrix $A\in\mathbb{R}^{n\times n}$ as
@@
Furthermore, we denote by $e_{i}$ the $i$th column of an identity matrix and by $\bm{1}$ a vector of ones, both of implicit length. The indicator function $\mathbbm{1}_{S}$ takes the value $1$ if the condition $S$ is true, and 00 otherwise.
@@ 2.1 Interacting-Langevin sampling
$\dot{U}_{t}=\nabla\log\widehat{\pi}(U_{t})+\sqrt{2}\,\dot{W}_{t}$ (6)
@@
$\dot{U}_{t}=K\nabla\log\widehat{\pi}(U_{t})+\sqrt{2K}\,\dot{W}_{t}.$ (7)
Denote by $\rho_{t}$ the law of $U_{t}$. Both Eq. 6 and Eq. 7 have the property that, under certain conditions on $\widehat{\pi}$, they transform arbitrary distributions $\rho_{0}$ of the initial value $U_{0}$ into a distribution $\rho_{\infty}\coloneqq\lim_{t\rightarrow\infty}\rho_{t}$ that is exactly the desired distribution $\pi$. This leads to a Monte Carlo algorithm for approximately sampling from $\pi$: simulate some number $J$ of discretized paths $\{U_{t}^{j}\}_{j=1}^{J}$ of Eq. 6 or Eq. 7 until some sufficiently large time $T$ and use $\{U_{T}^{j}\}_{j=1}^{J}$ as the approximate samples.
@@
$\dot{U}_{t}=\mathcal{C}(\rho_{t})\nabla\log\widehat{\pi}(U_{t})+\sqrt{2\mathcal{C}(\rho_{t})}\,\dot{W}_{t}.$ (8)
@@
Consider an SDE parametrized by the unnormalized target distribution $\widehat{\pi}$ (or, equivalently, by the potential $V$). The SDE is called affine-invariant if it is invariant under affine transformations of the state variables and $\widehat{\pi}$. That is, define
[…]
for an arbitrary invertible matrix $M$ and vector $b$. Denote by $\rho_{t}$ the probability density at time $t$ of solutions to the SDE parameterized by $\widehat{\pi}$ and with initial distribution $\rho_{0}$, and by $\tilde{\rho}_{t}$ the equivalent for $\tilde{\pi}$ and $\tilde{\rho}_{0}$. Then, the SDE is affine-invariant if
[…]
@@
Note that the SDE Eq. 8 is of McKean–Vlasov type [^1] These McKean–Vlasov equations can be simulated by evolving an ensemble of many time-discretized solution paths in parallel and using their empirical distribution instead of the law $\rho_{t}$. This is justified by the principle of propagation of chaos [56]; for the ensemble Kalman sampler, this was studied in detail in [14]. Until section 6, we study non-discretized McKean–Vlasov SDEs.: its drift and diffusion coefficients depend on the law $\rho_{t}$ through $\mathcal{C}(\rho_{t})$. It is also possible to use any positive definite preconditioner $\mathbf{C}(\rho_{t};U_{t})$ that depends on the current position $U_{t}$ as well as on its law $\rho_{t}$. Similarly to what was noted in [40, 47], the resulting SDE
$\dot{U}_{t}=\mathbf{C}(\rho_{t};U_{t})\nabla\log\widehat{\pi}(U_{t})+\sqrt{2\mathbf{C}(\rho_{t};U_{t})}\,\dot{W}_{t}$ (13)
has as its Fokker–Planck equation
@@
$\dot{U}_{t}=\mathbf{C}(\rho_{t};U_{t})\nabla\log\widehat{\pi}(U_{t})+\nabla_{u}\cdot\mathbf{C}(\rho_{t};U_{t})+\sqrt{2\mathbf{C}(\rho_{t};U_{t})}\,\dot{W}_{t}.$ (14)
@@ 2.2 Consensus-based sampling
Consensus-based sampling [6] is defined by the dynamics [^2] CBS also has an optimization mode where the factor $(\alpha+1)$ is omitted, which results in an optimization method instead of a sampling one. This paper only considers CBS in sampling mode.
$\dot{U}_{t}=-(U_{t}-\mu_{\alpha}(\rho_{t}))+\sqrt{2(\alpha+1)\,\mathcal{C}_{\alpha}(\rho_{t})}\,\dot{W}_{t}.$ (15)
@@ 3 Consensus-based sampling as approximate interacting Langevin
We now propose a new connection between interacting-Langevin and consensus-based samplers, arriving at the CBS dynamics Eq. 15 starting from Eq. 14 by introducing three specific approximations. Throughout the following motivation, we assume that the potential function $V$ is differentiable and $L$-weakly convex (meaning that $V(\cdot)+\frac{L}{2}\lVert\,\cdot\,\rVert^{2}$ is convex for some $L\in\mathbb{R}_{0}^{+}$). We also assume that $\mathcal{C}_{\alpha}(\rho_{t})$ is positive definite for all $t$.
@@
the $A$-Moreau envelope of $V$. It approximates the potential $V$, since $V^{A}(u)\nearrow V(u)$ for all $u$ as $\lVert A\rVert\searrow 0$ [50, §1G]. Let us take Eq. 14 with ${\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}_{\alpha}(\rho_{t})}$ as a starting point. We introduce scalar parameters $\kappa,\gamma>0$ and approximate $V$ in Eq. 14 by its $\gamma$-rescaled $\kappa\mathcal{C}_{\alpha}(\rho_{t})$-Moreau envelope $\gamma V^{\kappa\mathcal{C}_{\alpha}(\rho_{t})}$. We will later choose the parameter $\gamma$ to counteract the approximation error of the Moreau envelope. If $V$ is $L$-weakly convex and $\lVert A\rVert<L^{-1}$, then $V^{A}$ has gradient [50, Theorem 2.26]
$\nabla V^{A}(u)=(A)^{-1}(u-\operatorname{\mathbf{prox}}^{A}_{V}(u)),$ (19)
where the proximal operator denotes the singleton
@@
$\dot{U}_{t}=-\frac{\gamma}{\kappa}\left(U_{t}-\operatorname{\mathbf{prox}}^{\kappa\mathcal{C}_{\alpha}(\rho_{t})}_{V}(U_{t})\right)+\sqrt{2\mathcal{C}_{\alpha}(\rho_{t})}\,\dot{W}_{t}.$ (21)
@@
Solving a minimization problem to evaluate the proximal operator may be expensive. Instead note that the Laplace principle [12, 38] implies (similarly to, e.g., [5])
$\operatorname{\mathbf{prox}}^{\kappa\mathcal{C}_{\alpha}(\rho_{t})}_{V}(U_{t})=\lim_{\beta\rightarrow\infty}\mu_{\beta,\kappa\mathcal{C}_{\alpha}(\rho_{t})/\beta}(\rho_{t};U_{t})$ (22)
under mild conditions. Here, we used the weighted mean
$\mu_{a,A}(\rho;u)$ $\coloneqq\frac{\int vw_{a,A}(v;u)\,\mathrm{d}\rho(v)}{\int w_{a,A}(v;u)\,\mathrm{d}\rho(v)},$ (23)
$w_{a,A}(v;u)$ $\coloneqq\exp\bigl{(}-\lVert v-u\rVert_{A}^{2}/2\bigr{)}\widehat{\pi}(v)^{a}.$ (24)
@@
$\dot{U}_{t}=-\frac{\gamma}{\kappa}\Bigl{(}U_{t}-\mu_{\beta,\kappa\mathcal{C}_{\alpha}(\rho_{t})/\beta}(\rho_{t};U_{t})\Bigr{)}+\sqrt{2\mathcal{C}_{\alpha}(\rho_{t})}\,\dot{W}_{t}.$ (25)
@@
$\dot{U}_{t}=-(\alpha+1)^{-1}(U_{t}-\mu_{\alpha}(\rho_{t}))+\sqrt{2\mathcal{C}_{\alpha}(\rho_{t})}\,\dot{W}_{t}.$ (27)

### v3

@@ Abstract
We propose a localized consensus-based method for sampling from non-Gaussian distributions. This method arises from an alternative derivation of consensus-based sampling (CBS). Starting from ensemble-preconditioned Langevin dynamics, we approximate the potential with a Moreau envelope, replace the gradient in the Langevin equation with a proximal operator, and finally approximate this operator by a weighted mean. Under Gaussian initial and target distributions, this procedure recovers the standard CBS dynamics. In addition, when we retain only the approximations valid beyond the Gaussian case, we retrieve a refined variant of polarized CBS. The resulting algorithm, which we call localized consensus-based sampling, is affine-invariant, exact for Gaussian targets in the mean-field limit, and demonstrates improved robustness over polarized CBS in numerical experiments. Like other consensus-based methods, localized CBS is fully gradient-free and easily parallelizable.
@@ 1 Introduction
We consider the problem of sampling a $d$-dimensional probability distribution with density
@@
where $V\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a potential function with $\int\exp(-V(v))\mathrm{d}v<\infty$. Here, we implicitly take integrals over $\mathbb{R}^{d}$. In many applications $V$ is known but the integral $Z$ may be very expensive to compute, such that $\pi$ is available only up to a multiplicative constant.
@@ 1.1 Bayesian inverse problems
Equation 1 commonly arises in Bayesian inversion (see, e.g., [31, 56]). Consider a model $G$ with unknown parameters $u$ and perturbed observation
@@
where $\eta$ is some unknown measurement noise. The goal is characterizing likely values of $u$, based on $y$ and $G$. Denoting by $\pi_{\mathrm{prior}}(u)$ and $\pi_{\mathrm{noise}}(\eta)$ our prior beliefs about $u$ and $\eta$, Bayes’ rule describes our belief about $u$ after the observation by the posterior distribution with density
@@
Optimization algorithms can solve the inverse problem in this fashion. However, $u_{\mathrm{MAP}}$ gives no information about the uncertainty on $u$ captured by $\pi_{\mathrm{post}}$. When uncertainty quantification is required, other characteristics of the posterior must be inferred. In that case, one commonly wants to integrate some function with respect to $\pi_{\mathrm{post}}$, e.g., to compute statistical moments of the distribution. Numerical integration techniques, such as [26], which uses quasi-Monte Carlo, may provide a sufficient approximation even in high dimensions. Alternatively, it is common to use algorithms that produce a sequence of Monte Carlo samples distributed approximately according to $\pi_{\mathrm{post}}$. Then solving Eq. 2 corresponds to drawing samples from Eq. 1, with the evidence $\pi_{\mathrm{ev}}(y)$ as the intractable $Z$ and with $\widehat{\pi}_{\mathrm{post}}(\cdot\mid y)$ as the unnormalized density $\widehat{\pi}$.
@@ 1.2 Related literature
Sampling from distributions with density Eq. 1 with unknown $Z$ is historically done with Markov chain Monte Carlo (MCMC) algorithms, which originate in [25, 38] and are reviewed in detail in [3]. MCMC remains a very active research topic with recent developments such as the popular pCN [8] and DILI [9] proposal moves that are effective in high dimensions and multilevel MCMC versions [16, 36] to exploit cheap approximations to $G$.
An alternative method is sequential Monte Carlo (SMC) [11], which evolves samples from an easy-to-sample distribution (e.g., $\pi_{\mathrm{prior}}$ in Eq. 3) into samples from $\pi$ gradually over multiple steps. Many other methods are based on evolving SDEs that have $\pi$ (or an approximation to $\pi$) as an invariant distribution. An example is Langevin diffusion (see subsection 2.1); after time discretization, the unadjusted Langevin algorithm (ULA) [50] arises as an approximate sampler of $\pi$. Using ULA steps as MCMC proposals—or, equivalently, adding Metropolis–Hastings rejection to ULA—results in the Metropolis-adjusted Langevin algorithm (MALA) [2].
These Langevin methods use gradients of the potential $V$, which can be undesirable [17, 33]. Various methods adapt Langevin dynamics to be gradient-free, by using an ensemble of multiple particles (instances of the SDE) that exchange information, which can be harnessed to approximate gradients. These are often formulated in the infinite-particle limit, as McKean–Vlasov SDEs. Examples include ensemble Kalman sampling (EKS) [20] and the closely related ALDI method [21], derived from interacting Langevin diffusions and inspired by the ensemble Kalman inversion (EKI) optimizer [29, 33, 53]. EKI is related to iterating the ensemble Kalman filter (EnKF) [18]; an EKI-based sampler is analyzed in [14]. Other ensemble variants of Langevin dynamics include [13] and the multiscale sampler in [44]. In addition, the consensus-based sampling (CBS) method [6], described in subsection 2.2, was developed independently of Langevin dynamics. It was inspired by consensus-based optimization (CBO) [47].
Interaction in particle methods can be localized: instead of computing a global interaction term, each particle interacts more with particles that are nearby than with those that are far away. The aim of localization is to avoid the linearity assumption on $G$ that is implicit in these methods. Localized algorithms show improved performance for nonlinear $G$, since the adapted dynamics rely on local instead of global linearity. The paper [48] proposed to use localized interaction in the EKS/ALDI algorithm—details are given in subsection 5.3. Similar ideas were proposed for EKI and the EnKF [60], for an ensemble Kalman method for rare-event sampling [61], and for CBS as a variant named polarized CBS [4]—see subsection 5.2. Our paper develops a different CBS variant with localized interaction, named localized CBS.
@@
We stress that localization as discussed here is unrelated to the notion of localization in [27, 43] for the EnKF, in [1, 59] for EKI, and in [40] for MCMC. There, the empirical covariance matrix is corrected for spurious correlations, caused by a finite ensemble size, between parameter indices that should be uncorrelated—e.g., because they correspond to elements at a large distance in the underlying problem. This technique can also increase the rank of the covariance and prevent EKI’s subspace property [59].
@@ 1.3 Objectives and contributions
    Parallelizability. Standard MCMC methods are inherently sequential due to the need to overcome a burn-in phase and are hard to parallelize effectively [55, 62]. SMC and McKean–Vlasov-based methods use an ensemble of particles in parameter space that exchange information through interaction. This means that any work that is local to a single particle, such as evaluating $G$, can be trivially parallelized over all particles.
@@
    Non-Gaussian performance. Many sampling methods use approximations that hold only for Gaussian distributions, leading to poor sampling performance when the target distribution is non-Gaussian or multimodal.
@@
    Section 3 proposes a novel way to connect interacting Langevin diffusions [20], which use gradients, to consensus-based sampling [6], which does not. We rely on three approximations: replacing $V$ by a Moreau envelope, approximating the proximal operator by a weighted mean, and assuming that the target distribution is Gaussian.
@@
    Subsections 4.1 and 4.2 adapt these approximations, removing those that go against our objective of solving non-Gaussian sampling problems. This results in an algorithm we name localized consensus-based sampling. We perform an analysis of localized CBS with a broad class of covariance-based preconditioners. We derive ODEs for the evolution of its mean and covariance when applied to Gaussian problems in subsection 4.3. More generally, subsection 4.4 shows that localized CBS is affine-invariant.
@@
    Section 5 compares the existing CBS, polarized CBS, and localized ALDI algorithms to our localized CBS method. We argue that our algorithm is an improved version of polarized CBS [4], with our comparative benefits (such as affine-invariance and non-Gaussian performance) emerging from the connection to interacting Langevin diffusions.
@@
    Section 6 discusses approximations of the mean-field equations with finite particle systems and other computational considerations. Section 7 performs numerical experiments that showcase the method’s affine-invariance and non-Gaussian performance.
@@ 1.4 Notation
We denote by $e_{i}$ the $i$th column of an identity matrix and by $\bm{1}$ a vector of ones, both of implicit length. In addition, we define the Mahalanobis norm of a vector $v\in\mathbb{R}^{n}$ weighted by a positive definite matrix $A\in\mathbb{R}^{n\times n}$ as
@@
The indicator function $\mathbbm{1}_{S}$ takes the value $1$ if the condition $S$ is true, and 0 otherwise.
@@ 2.1 Interacting-Langevin sampling
$\mathrm{d}U_{t}=\nabla\log\widehat{\pi}(U_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}$ (6)
@@
$\mathrm{d}U_{t}=K\nabla\log\widehat{\pi}(U_{t})\,\mathrm{d}t+\sqrt{2K}\,\mathrm{d}W_{t}.$ (7)
Denote by $\rho_{t}$ the law of $U_{t}$. Both Eq. 6 and Eq. 7 have the property that, under certain conditions on $\widehat{\pi}$, they transform arbitrary laws $\rho_{0}$ of the initial value $U_{0}$ into a law $\rho_{\infty}\coloneqq\lim_{t\rightarrow\infty}\rho_{t}$} that is exactly the desired distribution $\pi$. This leads to a Monte Carlo algorithm for approximately sampling from $\pi$: simulate some number $J$ of discretized paths $\{U_{t}^{j}\}_{j=1}^{J}$ of Eq. 6 or Eq. 7 until some sufficiently large time $T$ and use $\{U_{T}^{j}\}_{j=1}^{J}$ as the approximate samples.
@@
$\mathrm{d}U_{t}=\mathcal{C}(\rho_{t})\nabla\log\widehat{\pi}(U_{t})\,\mathrm{d}t+\sqrt{2\mathcal{C}(\rho_{t})}\,\mathrm{d}W_{t}.$ (8)
@@
Consider an SDE that depends on the unnormalized target distribution $\widehat{\pi}$ (or, equivalently, on the potential $V$). The SDE is called affine-invariant if it is invariant under affine transformations of the state variables and $\widehat{\pi}$. That is, define
$\tilde{\pi}(z)=\widehat{\pi}(Mz+b)\qquad\text{and}\qquad\tilde{\rho}_{0}(z)=\mathinner{\!\left\lvert M\right\rvert}\rho_{0}(Mz+b)$ (11)
for an arbitrary invertible matrix $M$ and vector $b$. Denote by $\rho_{t}$ the probability density at time $t$ of solutions to the SDE for $\widehat{\pi}$ and with initial distribution $\rho_{0}$, and by $\tilde{\rho}_{t}$ the equivalent for $\tilde{\pi}$ and $\tilde{\rho}_{0}$. Then, the SDE is affine-invariant if
$\tilde{\rho}_{t}(z)=\mathinner{\!\left\lvert M\right\rvert}\rho_{t}(Mz+b)\qquad\text{for all $t$}.$ (12)
@@
Note that the SDE Eq. 8 is of McKean–Vlasov type [^1] These McKean–Vlasov equations can be simulated by evolving an ensemble of many time-discretized solution paths in parallel and using their empirical distribution instead of the law $\rho_{t}$. This is justified by the principle of propagation of chaos [57]; for the ensemble Kalman sampler, this was studied in detail in [15]. Until section 6, we study non-discretized McKean–Vlasov SDEs.: its drift and diffusion depend on the law $\rho_{t}$ through $\mathcal{C}(\rho_{t})$. It is also possible to use any positive definite preconditioner $\mathbf{C}(\rho_{t};U_{t})$ that depends on the state $U_{t}$ as well as on its law $\rho_{t}$. The resulting SDE is
$\mathrm{d}U_{t}=\mathbf{C}(\rho_{t};U_{t})\nabla\log\widehat{\pi}(U_{t})\,\mathrm{d}t+\sqrt{2\mathbf{C}(\rho_{t};U_{t})}\,\mathrm{d}W_{t}$ (13)
and, similarly to what was noted in [41, 48], has as its Fokker–Planck equation
@@
$\mathrm{d}U_{t}=\mathbf{C}(\rho_{t};U_{t})\nabla\log\widehat{\pi}(U_{t})\,\mathrm{d}t+\nabla_{u}\cdot\mathbf{C}(\rho_{t};U_{t})\,\mathrm{d}t+\sqrt{2\mathbf{C}(\rho_{t};U_{t})}\,\mathrm{d}W_{t}.$ (14)
@@ 2.2 Consensus-based sampling
Consensus-based sampling [6] follows the dynamics [^2] CBS also has an optimization mode where the factor $(\alpha+1)$ is omitted, which results in an optimization method instead of a sampling one. This paper only considers CBS in sampling mode.
$\mathrm{d}U_{t}=-(U_{t}-\mu_{\alpha}(\rho_{t}))\,\mathrm{d}t+\sqrt{2(\alpha+1)\,\mathcal{C}_{\alpha}(\rho_{t})}\,\mathrm{d}W_{t}.$ (15)
@@ 3 Consensus-based sampling as approximate interacting Langevin
We now propose a new connection between interacting-Langevin and consensus-based samplers, arriving at the CBS dynamics Eq. 15 starting from Eq. 14 by making three specific approximations. Throughout, we assume that the potential function $V$ is differentiable and $L$-weakly convex (meaning that $V(\cdot)+\frac{L}{2}\lVert\,\cdot\,\rVert^{2}$ is convex for some $L\in\mathbb{R}_{0}^{+}$) and that $\mathcal{C}_{\alpha}(\rho_{t})$ is positive definite for all $t$.
@@
the $A$-Moreau envelope of $V$. It approximates the potential $V$, since $V^{A}(u)\nearrow V(u)$ for all $u$ as $\lVert A\rVert\searrow 0$ [51, §1G]. Let us take Eq. 14 with $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}_{\alpha}(\rho_{t})$} as a starting point. We introduce scalar parameters $\kappa,\gamma>0$ and approximate $V$ in Eq. 14 by its $\gamma$-rescaled $\kappa\mathcal{C}_{\alpha}(\rho_{t})$-Moreau envelope $\gamma V^{\kappa\mathcal{C}_{\alpha}(\rho_{t})}$. We will later choose the parameter $\gamma$ to counteract the approximation error of the Moreau envelope. If $V$ is $L$-weakly convex and $\lVert A\rVert<L^{-1}$, then $V^{A}$ has gradient
$\nabla V^{A}(u)=(A)^{-1}(u-\operatorname{\mathbf{prox}}^{A}_{V}(u));$ (19)
see [51, Theorem 2.26]. Here, the proximal operator denotes the singleton
@@
$\mathrm{d}U_{t}=-\frac{\gamma}{\kappa}\left(U_{t}-\operatorname{\mathbf{prox}}^{\kappa\mathcal{C}_{\alpha}(\rho_{t})}_{V}(U_{t})\right)\,\mathrm{d}t+\sqrt{2\mathcal{C}_{\alpha}(\rho_{t})}\,\mathrm{d}W_{t}.$ (21)
Instead of a gradient, we now must evaluate the proximal operator.
@@
Under mild conditions
$\operatorname{\mathbf{prox}}^{\kappa\mathcal{C}_{\alpha}(\rho_{t})}_{V}(U_{t})=\lim_{\beta\rightarrow\infty}\mu_{\beta,\kappa\mathcal{C}_{\alpha}(\rho_{t})/\beta}(\rho_{t};U_{t}),$ (22)
as implied by the Laplace principle [12, 39] (similarly to, e.g., [5]). We used the weighted mean
$\mu_{a,A}(\rho;u)\coloneqq\frac{\int vw_{a,A}(v;u)\,\mathrm{d}\rho(v)}{\int w_{a,A}(v;u)\,\mathrm{d}\rho(v)},\qquad w_{a,A}(v;u)\coloneqq\exp\bigl(-\lVert v-u\rVert_{A}^{2}/2\bigr)\widehat{\pi}(v)^{a}.$ (23)
@@
$\mathrm{d}U_{t}=-\frac{\gamma}{\kappa}\Bigl(U_{t}-\mu_{\beta,\kappa\mathcal{C}_{\alpha}(\rho_{t})/\beta}(\rho_{t};U_{t})\Bigr)\,\mathrm{d}t+\sqrt{2\mathcal{C}_{\alpha}(\rho_{t})}\,\mathrm{d}W_{t}.$ (24)
Note that neither the proximal operator in Eq. 21 nor the weighted mean in Eq. 24 may seem inherently easier to compute than the gradient in Eq. 14. The eventual reason for these approximations is that weighted means are suitable to efficient gradient-free particle discretizations, such as the ones in [6] and section 6. A direct particle discretization of Eq. 14 would clearly require gradients. For now we will study the dynamics in McKean–Vlasov form, but one should keep in mind this motivation.
@@
$\mathrm{d}U_{t}=-(\alpha+1)^{-1}(U_{t}-\mu_{\alpha}(\rho_{t}))\,\mathrm{d}t+\sqrt{2\mathcal{C}_{\alpha}(\rho_{t})}\,\mathrm{d}W_{t}.$ (26)